OpenStudy (anonymous):

2. Given the function f(x) = log2(x + 6), find the value of f−1(3).

OpenStudy (anonymous):

@ranga

OpenStudy (xapproachesinfinity):

find the inverse first

OpenStudy (xapproachesinfinity):

y=log_2(x+6) ==> 2^y=x+6 ==> x=2^y-6 switch x and y y=2^x-6 so f^-1(x)=2^x-6 then evaluate